A company uses a job order cost system and applies manufacturing overhead costs to jobs using a predetermined overhead rate based on direct labour-hours.

The following data were extracted from the company's accounting records for the previous year:

Manufacturing overhead: estimated: $50,000 actual: $55,000

Direct labour hours: estimated: #20,000 hrs actual: #24,000 hrs

A job completed during the year incurred the following costs:

Direct Materials $4,000

Direct labour $1,500 (at $5 per DLH)


What was the under applied or over applied overhead for the year?
Answer #1

Overhead rate = Estimated overhead/Estimated labor hour

= 50000/20000

Overhead rate = $2.5 per labor hour

Overhead applied = 24000*2.5 = $60000

Actual overhead = $55000

Over applied overhead = 60000-55000 = $5000